# Onsager vortex clusters on a sphere

### v1

@@ V Microcanonical Monte Carlo Simulations
In this section, we show numerical results of Monte Carlo (MC) simulations of a large number of point vortices on a unit sphere. The simulations are performed for a range of increasing energies and at fixed angular momentum $L_{\zeta}=0$.
V.1 Sandwich state and the quadrupole moment
In our microcanonical MC simulations, we adopt the general scheme developed in Refs. Smith (1989); Smith and O’Neil (1990); Yu et al. (2016) which are based on the demon algorithm Creutz (1983). In particular, we closely follow the algorithm developed in Ref. Yu et al. (2016) for a neutral vortex point vortex system. We find that the sandwich state emerges as increasing energy. A comparison between the analytical predictions of the mean field theory and the microcanonical MC sampling for point vortices is shown in Fig. 8. The numerical results for the values of $\nu$, $Q_{0}$ and the upper bound of the quadrupole moment are in good agreement with the mean field predictions.
@@ V.1 Sandwich state and the quadrupole moment
Figure 8: A comparison between the mean field theory and MC simulations for $N=1000$ point vortices. (a) and (b) show results from the mean field theory and Monte Carlo sampling for the scaled vorticity $\tilde{\sigma}$, respectively. Here $\widetilde{\sigma}=\sigma$ for $E=0.005$, $E=0.01$, and $\widetilde{\sigma}=\sigma/20$ for $E=0.1$. The purpose for introducing the scaled vorticity $\tilde{\sigma}$ is to present the large energy range on a single color map. (c) The quadrupole moment $Q$ for increasing energy $E$ (circle red) and the mean field prediction Eq. (56) (solid blue). The quadrupole moment is well bounded by the maximum possible value $Q^{\rm max}_{s}$ predicted by the mean field theory. The inset in (c) shows the $N$-dependence behavior of $Q$ in the vicinity of the clustering transition. For finite $N$, the transition energy $E_{c}$ is finite and approaches to zero as $N\rightarrow\infty$. The finite values of $Q\sim 1/\sqrt{N}$ below the transition energy is nearly energy independent and is due to uncorrelated fluctuations. The data presented here are taken in runs of $40000$ MC steps at each value of energy.
V.2 Sandwich state vs. quadrupole state
Our theory predicts that the sandwich state has higher entropy than the quadrupole state (see Sec. IV.2), for the fixed values of energy, number of vortices, and zero angular momentum. In order to verify this analytical result numerically, we perform the MC sampling starting from a quadrupole configuration at a moderate energy. We find that the vortex distribution saturates at the sandwich state [Fig.9(a)(b)], indicating that indeed the sandwich state has greater statistical weight than the quadrupole state. This can be further verified by measuring directly the entropy of the initial and final distributions for a certain range of energies [Fig.9(c)]. Here the initial quadrupole distribution is obtained through sampling of the quadrupole mode $\sigma_{21}$ (or $\sigma_{22}$).
@@
Figure 9: (a) A point vortex system containing $N=1000$ vortices which is initially prepared at a quadrupole state with $E=0.06$ and $L_{\zeta}=0$. (b) The finial state after $160000$ MC steps and the final state is the sandwich state. (c) A comparison between the entropy of the initial quadrupole state and the final sandwich state. The sphere is divided into $40\times 40$ cells by longitude and latitude (these cells have different area). The entropy is calculated through Eq. (94). Here the entropy is negative because of the artificial choice of the minimum area in the phase space, and only the entropy difference between the two states has physical meaning.
The entropy of a point vortex system confined on a domain can be measured numerically. We divide the domain into a large number of cells with size $a_{i}$, where $i$ denotes the cell index. Each cell contains a sufficient number of vortices. We denote the number of point vortices in a cell of each species as $N_{i,\hat{\sigma}}$ with $\hat{\sigma}=\pm$. The number of states is then given by
$\mathcal{W}\left(\{N_{i,\hat{\sigma}}\}\right)=\prod_{\hat{\sigma}}\left\{\prod_{i}\frac{N_{\hat{\sigma}}!}{N_{i,\hat{\sigma}}!}\left(\frac{a_{i}}{\Omega}\right)^{N_{i,\hat{\sigma}}}\right\},$ (94)
where is $\Omega$ is the area of domain Joyce and Montgomery (1973); Maestrini and Salman (2019). As it follows the entropy per vortex is $S=(1/N)\log\mathcal{W}$. We apply this method to vortices on a unit sphere and compare the entropy between the sandwich state and the quadrupole state [Fig.9(c)].
VI Conclusion and discussion
@@
We investigate Onsager vortex clustered states of point vortices confined on a sphere. We find that for vortex distributions with zero angular momentum, the sandwich state is the most probable state and is characterized by the quadrupole moment tensor. The mean field analytical predictions agree well with the Monte Carlo simulations. It is worthwhile mentioning that Ref. Kanai and Guo (2021) shows that the Gross-Pitaevskii dynamics drives an initial uniform vortex state to the end quadruple state through the evaporative heating process. This result is not necessary inconsistent with the our findings. Since the quadruple distribution also solves the self-consistent equation and hence it is a local maximum entropy state. In the situation described in Ref. Kanai and Guo (2021), vortices may be dynamically trapped in a local equilibrium state. To explore the entropy barrier between the sandwich state and the quadruple state and its effect on the dynamical transitions between the two states deserves future investigations.
@@ Acknowledgement
We thank Y. Xiong and C. Ma for useful discussions. X.Y. acknowledges support from the National Natural Science Foundation of China (Grant No. 12175215), the National Key Research and Development Program of China (Grant No. 2022YFA 1405300) and NSAF (Grant No. U2330401).
Appendix A Appendix
Appendix B Normolization dependence of the onset of the supercondensation
The value of $\beta$ at which the supercondensation occurs depends on the normalization condition. Here we show the method we used to determine the value of $\beta_{s}$ for the sandwich state [see Fig. 10(a)].
Figure 10: (a) The $\beta-e^{\gamma_{+}}$ curve is determined by normalization Eq. (73) and touches the $\beta=\beta_{s}=-16\pi$ curve when the supercondensation occurs. (b) The $\beta-e^{\gamma_{-}}$ curve is determined by the normalization Eq. (73) ($n_{+}\rightarrow n_{-}$). The condition $\beta e^{\gamma_{-}}<-5/2$ ensures $n^{\prime}_{-}(\rho)|_{\rho=0}=0$.

### v2

@@ IV Monte Carlo Simulations
In this section, we show numerical results of Monte Carlo (MC) simulations of a large number of point vortices on a unit sphere. The simulations are performed for a range of increasing energies and at fixed angular momentum $L_{\zeta}=0$ for the sandwich and the quadrupole state and $L_{\zeta}\neq 0$ for the dipole state.
IV.1 Microcanonical Monte Carlo
In our microcanonical Monte Carlo (MC) simulations, we adopt the general scheme developed in Refs. Creutz (1983); Smith (1989); Smith and O’Neil (1990); Yu et al. (2016), where the total energy $E$ and the angular momentum $L_{\zeta}$ are fixed within narrow shells. For vortices on a unit sphere, $q_{i}=\cos\theta_{i}$ and $p_{i}=\phi_{i}$ are canonical variables Kimura (1999). In order to keep sampling vortices uniformly in the phase space, at each Monte Carlo step, we randomly choose a vortex dipole at positions $\bm{r}^{+}(\theta_{i},\phi_{i})$ and $\bm{r}^{-}(\theta_{i},\phi_{i})$ and move it to $\bm{r}^{+}(\theta^{\prime}_{i},\phi^{\prime}_{i})$ and $\bm{r}^{-}(\theta^{\prime}_{i},\phi^{\prime}_{i})$, where $\phi^{\prime}=p$, $\theta^{\prime}_{i}=\arccos(q)$ with $p\in[0,2\pi]$ and $q\in[-1,1]$being independent and uniformly distributed random variables. The new configuration is accepted if the updated energy $E^{\prime}$ and the updated angular momentum $L^{\prime}_{\zeta}$ are within the narrow shells. Otherwise, the vortex configuration remains the same as the previous one.
IV.2 The sandwich state vs. the quadrupole state
By appropriately choosing initial vortex distributions for given energies and zero angular momentum, the sandwich state appears at the end of our MC sampling. A comparison between the analytical predictions and the numerical results is shown in Fig. 9 and excellent agreements are found. Caution should be taken here, it is that the initial configures are chosen to be not far from the sandwich state due to our limited MC sampling efficiency. It is difficult for our current MC sampling method to select the sandwich state but not the quadrupole state from a random initial vortex distribution within the energy and angular momentum constraints.
@@ IV.1 Microcanonical Monte Carlo
Figure 9: A comparison between the mean field theory and MC simulations for $N=1000$ point vortices. (a) and (b) show results from the mean field theory and Monte Carlo sampling for the scaled vorticity $\tilde{\sigma}$, respectively. Here $\widetilde{\sigma}=\sigma$ for $E=0.005$, $E=0.01$, and $\widetilde{\sigma}=\sigma/10$ for $E=0.1$. The purpose for introducing the scaled vorticity $\tilde{\sigma}$ is to present the large energy range on a single color map. 0 (c) The quadrupole moment $Q_{s}$ for increasing energy $E$ (circle red) and the mean field prediction Eq. (61) (solid blue). The quadrupole moment is well bounded by the maximum possible value $Q^{\rm max}_{s}$ predicted by the mean field theory. The inset in (c) shows the NN-dependence behavior of $Q_{s}$ in the vicinity of the clustering transition. For finite $N$, the transition energy $E_{c}$ is finite and approaches zero as $N\rightarrow\infty$. The finite values of $Q_{s}\sim 1/\sqrt{N}$ below the transition energy are nearly energy independent due to uncorrelated fluctuations. The data presented here are taken in runs of 4000040000 MC steps at each value of energy.
To verify that the sandwich state is the maximum entropy state for large but finite $N$, we numerically evaluate the entropy values of the sandwich state and the quadrupole state at a certain energy range. The entropy values are evaluated using Eq. (17). We uniformly divide the phase space of a vortex on a unit sphere into a large number of cells labeled by $q_{i},p_{j}]$ with $q_{i}=-1+2(i-1)/\tilde{N}$, $p_{j}=2\pi(j-1)/\tilde{N}$ and $j=1,...\tilde{N}+1$, where $(i,j)$ denotes the cell index and $\tilde{N}^{2}$ is number of the cells. The area of a cell is then $a_{i}=4\pi/\tilde{N}^{2}$. In terms of spherical coordinates, the cell labels are $\theta_{i},\phi_{i}]$ with $\theta_{i}=\arccos q_{i}$ and $\phi_{j}=p_{i}$. Note that here $\tilde{N}$ should be properly chosen such that most cells contain a sufficient large number of vortices. Using this method we confirm that the sandwich state has higher entropy than the quadrupole state (Fig.10).
@@
Figure 10: A comparison between the entropy of the quadrupole state and the sandwich state. Here $\tilde{N}=32$ and the quadrupole and the sandwich states are prepared via MC sampling for $N=4000$ point vortices with initial vortex distributions very closed to the quadrupole state and the sandwich state, respectively. The entropy is negative because of the artificial choice of the minimum area in the phase space, and only the entropy difference between the two states has physical meaning.
IV.3 The dipole state
Our perturbation analysis at low energies (Sec. III.3.1) shows that for $L_{\zeta}\neq 0$ and $\omega=0$ the dipole state maximizes the entropy for a neutral vortex system on a sphere. This is different from neutral vortices confined to a disc where the dipole state carries zero angular momentum Yu et al. (2016). In this subsection, we present the MC simulations and compare them with our analytical predictions, showing excellent agreements (Fig.11).
Figure 11: A comparison between the analytical prediction $D=|D_{\zeta}|=4\sqrt{\pi/3}\sqrt{E}$ [Eq. (51) ](solid black) and the MC sampling for $N=1000$ point vortices (circles) in the dipole state. The dashed blue line marks the upper bound of the dipole moment $D^{\rm max}=2$ which is reached at the supercondensate point. Note that the angular momentum $L_{\zeta}=4\sqrt{\pi/3}\sqrt{E}$ also increases as increasing energy for this rotation free ($\omega=0$) dipole state.
V Conclusion and discussion
@@
We investigate rotation free Onsager vortex clustered states of point vortices confined on a sphere. We find that the sandwich state is the most probable state with zero angular momentum and is characterized by the quadrupole moment tensor. The dipole state is the maximum entropy state with finite angular momentum and is characterized by nonzero dipole moment. Our mean field analytical predictions agree well with the Monte Carlo simulations. In Ref. Kanai and Guo (2021), it is shown that the point-vortex dynamics combined removing low energy vortex pairs drives an initial uniform vortex state to the quadruple state. This result is not necessarily inconsistent with our findings, since the quadruple distribution also solves the self-consistent equation and hence is a local maximum entropy state. In the situation described in Ref. Kanai and Guo (2021), vortices may be dynamically trapped in a local equilibrium state. To explore the entropy barrier between the sandwich state and the quadruple state and its effect on the dynamical transitions between the two states deserves future investigations.
@@ Acknowledgement
We thank C. Ma and Y. Xiong for useful discussions. X.Y. acknowledges support from the National Natural Science Foundation of China (Grant No. 12175215, Grant No. 12475041), the National Key Research and Development Program of China (Grant No. 2022YFA 1405300) and NSAF (Grant No. U2330401).
Appendix A Comparison of entropy between $\ell=2$ and $\ell>2$ modes
To leading order in $\delta E$, the entropy for different modes is obtained in the main text and reads
[…]
where $\beta_{c,\ell}=-\ell(\ell+1)/(2n_{0})$. Let us consider $\ell=2,m=0$ and $\ell=3,m=0$ states:
$\psi_{20}=$ $c_{20}{\rm Y}_{20}=\sqrt{\frac{E_{0}}{3}}\frac{1}{4}\sqrt{\frac{5}{\pi}}(3\cos^{2}\theta-1),$ (93)
$\psi_{30}=$ $c_{30}{\rm Y}_{30}=\sqrt{\frac{E_{0}}{6}}\frac{1}{4}\sqrt{\frac{7}{\pi}}\left(5\cos^{3}\theta-3\cos\theta\right),$ (94)
and evaluate the corresponding values of the entropy using Eq. (92). We find that
[…]
and clearly $S[{\ell=2,m=0}]>S[{\ell=3,m=0}]$. In general, since $\beta^{2}_{c,\ell>2}-\beta^{2}_{c,\ell=2}\sim[(\ell-2)^{2}+5(\ell-2)]^{2}$ and $\int{\rm d}\Omega\,\psi^{2}_{\ell m}\sim\mathcal{O}{(1)}$, we can conclude that $\ell=2$ have larger entropy than the modes with $\ell>2$.
Appendix B Boundary conditions
In the main text, we report that at high energies the vortex distributions near the poles and the equator (in the stereographic coordinates) can be well-approximated by exact solutions to Eq. (68) [or Eq (75)] describing distributions of chiral vortices confined to a disc with proper boundary conditions. Here we show that these boundary conditions are induced be the flow generated by the vortices on the sphere and can be obtained from the global stream function Eq. (54) at low energies.
The relevant stream function Eq. (54) is
and hence the velocity field reads
$u_{\theta}=\frac{1}{\sin\theta}\frac{\partial\psi}{\partial\phi},\quad u_{\phi}=-\frac{\partial\psi}{\partial\theta}.$ (98)
It is clear that
$u_{\phi}(\theta=0)=u_{\phi}(\theta=\pi)=u_{\phi}(\theta=\pi/2)=0.$ (99)
By the stereographic projection $\rho=\cot(\theta/2)$], Eq (97) becomes
$\psi_{20}(\rho)=\frac{c_{20}}{4}\sqrt{\frac{5}{\pi}}\left(3\left(\frac{\rho^{2}-1}{\rho^{2}+1}\right)^{2}-1\right)$ (100)
and the velocity field (99) induces that
$\psi^{\prime}(\rho)|_{\rho=0}=\psi^{\prime}(\rho)|_{\rho=1}=0$ (101)
which are the adopted boundary conditions in the main text.
Let us now discuss the irrelevance of the values of $\psi(\rho=0)$ and $\psi(\rho=1)$. If applying the boundary condition $\psi^{\prime}(\rho)|_{\rho=0}=0$ while without specifying the value of $\psi(\rho=0)$, the solution to Eq. (69) reads
$\psi(\rho)=-\frac{2}{\beta}\log\left(\frac{2\sqrt{C}}{\beta\rho^{2}{\rm e}^{\gamma_{+}}C-2}\right)$ (102)
and the vortex density reads
$n_{+}(\rho)=\exp(-\beta\psi+\gamma_{+})=\frac{4{\rm e}^{\gamma_{+}}C}{(\beta\rho^{2}{\rm e}^{\gamma_{+}}C-2)^{2}},$ (103)
where $C$ is an arbitrary constant. Clearly here the overall factor ${\rm e}^{\gamma_{+}}C$ can be determined by the normalization condition
$\frac{1}{2}=\int^{2\pi}_{0}{\rm d}\phi\int_{0}^{1}\frac{4\rho}{(1+\rho^{2})^{2}}\,{\rm d}\rho\,n_{+}(\rho),$ (104)
and the value of $C$ is irrelevant. In the main text, we choose $\psi(\rho=0)=0$, which gives rise to $C=1$, for convenience.
Similarly, if applying the boundary condition $\psi^{\prime}(\rho)|_{\rho=1}=0$ while without specifying the value of $\psi(\rho=1)$, the solution to Eq. (76) reads
$\psi(\rho)=\frac{1}{\beta}\log\left(\frac{A^{2}(4-A^{2})\rho^{(-2+A)}}{2\beta{\rm e}^{\gamma_{-}}(A+2-(A-2)\rho^{A})^{2}}\right)$ (105)
and the vortex density is
$n_{-}(\rho)=\frac{2A^{2}(4-A^{2})\rho^{A-2}}{\beta\left(A\rho^{A}-2\rho^{A}+A+2\right)^{2}},$ (106)
where $A$ is a constant and can be determined by the normalization condition
$\frac{1}{2}=\int^{2\pi}_{0}{\rm d}\phi\int_{0}^{1}\frac{4\rho}{(1+\rho^{2})^{2}}\,{\rm d}\rho\,n_{-}(\rho).$ (107)
Hence the value of $\psi(\rho=1)$ is irrelevant and only specifies the relation between $A$ and $\gamma_{-}$. With the choice $\psi(\rho=1)=0$ which we made in the main text, we have $A=\sqrt{4-2\beta{\rm e}^{\gamma_{-}}}$ and hence the normalization condition determines $\gamma_{-}$.
Appendix C Normalization dependence of the onset of the supercondensation
The value of $\beta$ at which the supercondensation occurs depends on the normalization condition. Here we show the method we used to determine the value of $\beta_{s}$ for the sandwich state [see Fig. C(a)].
Figure C: (a) The $\beta-{\rm e}^{\gamma_{+}}$ curve is determined by normalization Eq. (72) and touches the $\beta=\beta_{s}=-16\pi$ curve when the supercondensation occurs. (b) The $\beta-{\rm e}^{\gamma_{-}}$ curve is determined by the normalization Eq. (72) ($n_{+}\rightarrow n_{-}$). The condition $\beta{\rm e}^{\gamma_{-}}<-5/2$ ensures $n^{\prime}_{-}(\rho)|_{\rho=0}=0$.